The length of a rectangle is four less than three times the width. The area is 20 square centimeters. Find the dimensions of the rectangle.
Answer by checkley75(3666)

L=3W-4
W(3W-4)=20
3W^2-4W-20=0
(3W-10)(W+2)=0
3W-10=0
3W=10
W=10/3 OR 3.333 ANSWER
L=3*3.333-4
L=10-4
L=6 ANSWER.
PROOF
3.333*6=20
20=20